OpenStudy (anonymous):

can you correct my answers? The first derivative of -t^2(5t+2)^3 is -2t(5t+2)^3 - 15t^2(5t+2)^2 The first derivative of sin^2(3t^2) is 12t(sin 3t^2)(cos 3t^2) and The first derivative of -5t^2 cos (2t^3) is -10t (cos 2t^3) + 30t^4 sin 2t^3

ganeshie8 (ganeshie8):

frist one is correct

ganeshie8 (ganeshie8):

[ sin^2(3t^2)]' = 2sin(3t^2) [(sin(3t^2)]' = 2sin(3t^2) cos(3t^2) [3t^2]' = 2sin(3t^2) cos(3t^2) 6t second one is also correct

OpenStudy (anonymous):

how about the last one?

OpenStudy (loser66):

correct, too

OpenStudy (anonymous):

many thanks @Loser66 :D

OpenStudy (anonymous):

thanks a lot too @ganeshie8 :)

OpenStudy (loser66):

ok